# Quantum Advantage in Distributed Sensing with Noisy Quantum Networks

### v1

@@ Abstract
We see that in the asymptotic regime of large $d$, the above fidelity threshold reduces to $1/d$. Moreover, the quantum advantage in DQS has close relationship with quantum entanglement, and the entanglement properties of the depolarized GHZ state have been well studied: $d$-qubit depolarized GHZ states are not completely separable if $F>3/(2^{d}+2)$ [61], and genuine multipartite entangled (GME) if $F>1/2$ [62]. It can be shown that $F_{\mathrm{th,dp}}>3/(2^{d}+2)$ for $d\geq 2$, which means that entanglement is necessary for DQS quantum advantage. Meanwhile, except for the special case of $\left.F_{\mathrm{th,dp}}\right|_{d=3}\approx 0.50963>1/2$, $F_{\mathrm{th,dp}}<1/2$ for $d>3$, which suggests that GME is in principle unnecessary to demonstrate quantum advantage in DQS. Further details can be found in Sec. III.A of the Supplemental Material [59].
@@
Quantum network simulation.—Following the above analytics, we simulate GHZ state distribution given realistic first-generation quantum repeater network [66, 67, 68] protocol stacks with an open-source, customizable quantum network simulator, SeQUeNCe [69]. We consider a simple 3-node network topology, where a center node directly connects the other two end nodes through optical fibers, with a Bell state measurement station in the middle of each fiber link. First, bipartite entanglement links are estabilished between the center node and the other nodes. Then LOCC, such as gate teleportation [70, 71, 72] and graph state fusion [73, 74], are able to generate a GHZ state across all network nodes.
@@
We have primarily focused on probe state preparation errors, while the parameter encoding dynamics can also be noisy due to sensor decoherence. It could be interesting to evaluate the impact of sensor decoherence after state preparation [81], where the time dependence of estimation accuracy gain [82, 83, 84] becomes important. We note that quantum error correction [85, 86, 87, 88] offers a good opportunity in fighting against noise in sensing process. Meanwhile, protocols such as continuous entanglement distribution [89, 90, 91] might be necessary to reduce latency of probe state preparation over quantum networks. In addition, optimization of Bell-state-based graph state distribution [92, 93, 94, 95, 96, 97, 98, 99, 100] is important for larger scale DQS. Finally, although we mainly assume finite-dimensional matter-based quantum sensors, photonic systems [101, 102, 103, 104, 105, 106] are also playing an important role in DQS.

### v2

@@ Abstract
We see that in the asymptotic regime of large $d$, the above fidelity threshold reduces to $1/d$. Moreover, the quantum advantage in DQS has a close relationship with quantum entanglement, and the entanglement properties of the depolarized GHZ state have been well studied: $d$-qubit depolarized GHZ states are not completely separable if $F>3/(2^{d}+2)$ [61], and genuine multipartite entangled (GME) if $F>1/2$ [62]. It can be shown that $F_{\mathrm{th,dp}}>3/(2^{d}+2)$ for $d\geq 2$, which means that entanglement is necessary for DQS quantum advantage. Meanwhile, except for the special case of $\left.F_{\mathrm{th,dp}}\right|_{d=3}\approx 0.50963>1/2$, $F_{\mathrm{th,dp}}<1/2$ for $d>3$, which suggests that GME is in principle unnecessary to demonstrate quantum advantage in DQS. Further details can be found in Sec. III.A of the Supplemental Material [59].
@@
Quantum network simulation.—Following the above analytics, we simulate GHZ state distribution given realistic first-generation quantum repeater [66, 67, 68] protocol stacks with an open-source, customizable quantum network simulator, SeQUeNCe [69]. We consider a simple 3-node network with linear topology, where a center node directly connects the other two end nodes through optical fibers, with a Bell state measurement station in the middle of each fiber link. First, bipartite entanglement links are established between the center node and the other nodes. Then LOCC, such as gate teleportation [70, 71, 72] and graph state fusion [73, 74], are performed to generate a GHZ state across all network nodes.
@@
We have primarily focused on probe state preparation errors, while the parameter encoding dynamics can also be noisy due to sensor decoherence. It could be interesting to evaluate the impact of sensor decoherence after state preparation [81], where the time dependence of estimation accuracy gain [82, 83, 84] becomes important. We note that quantum error correction [85, 86, 87, 88] offers a good opportunity in fighting against noise in sensing process. Meanwhile, protocols such as continuous entanglement distribution [89, 90, 91] might be necessary to reduce latency of probe state preparation over quantum networks. In addition, optimization of Bell-state-based graph state distribution [92, 93, 94, 95, 96, 97, 98, 99, 100] is important for larger scale DQS. Notably, privacy and security [101, 102, 103] is a potentially important aspect of realistic quantum sensor network as well. Finally, although we mainly assume finite-dimensional matter-based quantum sensors, photonic systems [104, 105, 106, 107, 108, 109] are also playing an important role in DQS.